OpenStudy (anonymous):

5(a-4)-8a=55 I will give medal to whoever help me thanks :)

OpenStudy (anonymous):

a= -25 distribute the 5 combine like terms sole for a

OpenStudy (anonymous):

5(a-4)-8a=55 5a-20-8a=55 -3a-20=55 -3a=75 a=-25